# Risk-Calibrated Affective Speech Recognition via Conformal Coverage Guarantees: A Stochastic Calibrative Framework for Emergent Uncertainty Quantification

### v1

@@ 4.1.1 Datasets
We used the Interactive Emotional Binary Motion Capture Database (IEMOCAP) and the Emotional Speech Recognition Dataset (TESS) for model training and evaluation. IEMOCAP contains about 12 hours of audio - visual data, including video, speech, facial motion capture, and text transcription. The data is performed by 10 actors (5 men and 5 women) in 5 sessions, covering emotion categories such as happiness, anger, sadness, frustration, and neutral. TESS is recorded by two actresses (26 and 64 years old, respectively), using the carrier phrase ”Say the word _” to express 200 target words, covering seven emotions: anger, disgust, fear, happiness, surprise, sadness, and neutral. There are a total of 2,800 audio files in WAV format.
(a) ResNet50 training and testing Loss-Acc
@@
4.1.2 Base Models
In this study, we used several convolutional neural network (CNN) architectures, including basic CNN, ResNet50, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet. We ensured that the number of parameters of these models was similar: ResNet50 (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters). These models optimize computational efficiency and performance through different designs. ResNet50 uses residual connections to solve the gradient vanishing problem of deep networks, MobileNetV3 and ShuffleNetV2 use lightweight designs to adapt to resource - constrained environments, and SqueezeNet and GhostNet further optimize the number of model parameters and computational complexity, making them suitable for embedded and mobile device applications.
4.1.3 Methods for Speech Feature Extraction
4.1.4 Evaluation Metrics
4.1.5 Hyper - parameters
We set the split ratio between the calibration set and the test set to 50% in the IEMOCAP dataset and 50% in the TESS dataset. We modified the structure of the models to ensure that each model has a similar number of parameters: ResNet50 (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters) to evaluate the uncertainty of different models with the same parameter size.
4.2 Determination of Speech Feature Extraction Methods
In the pre - experimental stage, we compared different speech feature extraction methods, including mel - spectrogram features, chromaticity features, and contrast features. We tested the TESS dataset and found that the accuracy of the mel - spectrogram feature on the test set reached 99.82%, significantly better than the chromaticity feature (97.86%) and contrast feature (87.41%). Based on these results, the mel - spectrogram feature showed stronger recognition ability and robustness, so we chose the mel - spectrogram as the optimal feature extraction method and widely used it in the training and evaluation of all models in subsequent experiments.
@@ 4.2 Determination of Speech Feature Extraction Methods
We trained five models (ResNet50, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet), all with a similar number of parameters. During training, mel - spectrogram features were used and trained on the IEMOCAP dataset (60% training set, 40% test set). All models achieved nearly 100% accuracy on the training set, but performed differently on the test set. Specifically, the test set accuracy of ResNet50 was 44.75%, MobileNetV3 was 40.56%, ShuffleNetV2 was 42.38%, SqueezeNet was 37.96%, and GhostNet was 35.06%. These results show that although the models performed well on the training set, they had weak generalization ability on the test set, reflecting the overfitting phenomenon, which provides a basis for the subsequent use of the Conformal Prediction framework for model evaluation.
4.3 Empirical Results of the ECR metric
We used SCP and Risk Control frameworks to evaluate the coverage performance of five models on the IEMOCAP dataset, and applied these models to the TESS dataset for testing. On the IEMOCAP and TESS datasets, both frameworks provided reliable coverage guarantees and showed that the two frameworks had good generalization capabilities across different datasets.
@@ 4.4 Uncertainty Estimation of the Model through the APSS metric
$0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.7$ $0.8$ $0.9$ $1$
IEMOCAP ResNet50 $3.757$ $2.856$ $2.186$ $1.619$ $1.176$ $0.839$ $0.560$ $0.318$ $0.139$ $0.025$
MobileNetV3 $4.158$ $3.132$ $2.440$ $1.847$ $1.369$ $0.982$ $0.645$ $0.380$ $0.155$ $0.066$
ShuffleNetV2 $3.964$ $2.991$ $2.306$ $1.738$ $1.310$ $0.914$ $0.597$ $0.345$ $0.151$ $0.050$
SqueezeNet $4.110$ $3.235$ $2.490$ $1.937$ $1.463$ $1.061$ $0.699$ $0.416$ $0.182$ $0.004$
GhostNet $4.357$ $3.457$ $2.752$ $2.169$ $1.650$ $1.214$ $0.808$ $0.456$ $0.200$ $0.011$
TESS ResNet50 $6.000$ $6.000$ $6.000$ $6.000$ $3.152$ $2.324$ $1.601$ $0.767$ $0.312$ $0.234$
MobileNetV3 $6.000$ $6.000$ $6.000$ $3.832$ $3.346$ $2.877$ $2.333$ $1.590$ $0.619$ $0.100$
ShuffleNetV2 $6.000$ $6.000$ $6.000$ $3.578$ $2.875$ $2.104$ $1.308$ $0.576$ $0.192$ $0.140$
SqueezeNet $6.000$ $6.000$ $6.000$ $5.979$ $3.490$ $2.855$ $2.090$ $0.878$ $0.299$ $0.112$
GhostNet $6.000$ $6.000$ $6.000$ $6.000$ $5.892$ $3.385$ $2.685$ $1.524$ $0.497$ $0.031$
Table 1: APSS under different data sets and risk levels
@@
We found a negative correlation between the average prediction set size (APSS) and the risk level. Specifically, as the risk level increases, the APSS value decreases significantly, indicating that the model’s prediction results at high risk levels are more certain and the size of the prediction set is optimized. In this way, APSS is able to effectively evaluate the uncertainty of the classification model on the test set. Larger APSS values indicate that the model has higher prediction uncertainty for certain categories, while smaller APSS values mean that the model has higher confidence in its predictions. Therefore, APSS, as a measure of model uncertainty, demonstrates its potential as a promising benchmark in evaluating the robustness of classification models and provides a valuable reference for further optimizing models.
@@ 4.5.1 Various Split Ratios
The experimental results show that although the calibration set size changes significantly (from 10% to 90%), this method always ensures the empirical coverage of the test set satisfies the theoretical lower bound $1-\alpha$ ($\alpha=0.2$) by dynamically adjusting the confidence threshold $q$. Theoretical analysis and empirical results jointly prove that this method exhibits strong robustness to data partition ratios. The constructed prediction sets provide statistically rigorous reliability guarantees for speech emotion recognition tasks, effectively addressing the confidence calibration failure problem in traditional methods caused by data distribution shifts.
@@ 5 Conclusion
This study addresses the issue of unreliable predictions in traditional neural networks for speech emotion recognition by proposing a solution based on CP and a Risk Control framework. The results indicate that traditional point prediction methods exhibit significant confidence bias due to factors such as data distribution differences, model overfitting, and difficulty in capturing complex emotional features. By using the SCP framework and the risk control framework, this study successfully transforms single-point predictions into set predictions with coverage guarantees, ensuring that, at a specified risk level $\alpha$, the true label is included in the prediction set with at least a probability of $1-\alpha$. Additionally, cross-dataset validation (where a model trained on the IEMOCAP dataset still maintains coverage guarantees on the TESS dataset) demonstrates the robustness of the framework under different data distributions.
We found a significant negative correlation between APSS and risk level, which provides an actionable metric for quantifying model decision uncertainty. Cross-dataset validation results show that even on the TESS dataset, with a large distribution difference, this metric still effectively reflects changes in model confidence. By constructing a dynamic relationship model between the prediction set size and risk level, this study establishes the first uncertainty evaluation standard suitable for speech emotion recognition.
This study adopts a set prediction method based on inconsistency score, and achieves strict coverage guarantee by dynamically adjusting the prediction set threshold. In addition, a more flexible risk control framework is introduced to balance the prediction set size and coverage probability by improving the loss function, thereby optimizing decision efficiency while ensuring statistical reliability. This framework can also be applied to problems with different needs by modifying the definition of the loss function. The study also found that APSS is negatively correlated with the risk level, indicating that when the risk level is high, the model’s prediction set is usually smaller and the model’s prediction uncertainty is lower. Based on this finding, a metric for evaluating the uncertainty of classification models on the test set is proposed.

### v2

@@ 4.1.1 Datasets
This study conducts model training and evaluation based on two multimodal emotion datasets, IEMOCAP and TESS. The IEMOCAP dataset contains 12 hours of audio-visual data collected across 5 recording sessions, where ten professional actors (5 male, 5 female) simulate real-life interaction scenarios.Busso et al. (2008)For model training, six emotion categories are selected: "ang" (anger), "hap" (happiness), "exc" (excitement), "sad" (sadness), "fru" (frustration), and "neu" (neutral). A stratified sampling approach is employed to ensure each category provides 2,000 balanced training samples. The TESS dataset comprises 2,800 audio files, in which two female actors express seven discrete emotions through the fixed carrier sentence structure "Say the word_".Pichora-Fuller and Dupuis (2020)To establish cross-dataset consistency, only emotion categories overlapping with IEMOCAP are retained. Specifically, "Anger", "Happiness", "Sadness", and "Neutral" in TESS are mapped to the corresponding IEMOCAP labels "ang", "hap", "sad", and "neu", respectively. This label harmonization ensures uniform emotional category definitions during model training and evaluation across datasets.
4.1.2 Base Models
In this study, we used several convolutional neural network (CNN) architectures, including basic CNN, ResNet, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet. We ensured that the number of parameters of these models was similar: ResNet (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters). These models optimize computational efficiency and performance through different designs. ResNet50 uses residual connections to solve the gradient vanishing problem of deep networksHe et al. (2016), MobileNetV3 and ShuffleNetV2 use lightweight designs to adapt to resource - constrained environmentsHoward et al. (2019); Ma et al. (2018), and SqueezeNet and GhostNet further optimize the number of model parameters and computational complexity, making them suitable for embedded and mobile device applications.Iandola et al. (2016); Han et al. (2020)
4.1.3 Methods for Speech Feature Extraction
4.1.4 Evaluation Metrics
4.1.5 Hyper - parameters
We set the split ratio between the calibration set and the test set to 50% in the IEMOCAP dataset and 50% in the TESS dataset. We modified the structure of the models to ensure that each model has a similar number of parameters: ResNet (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters) to evaluate the uncertainty of different models with the same parameter size.
4.2 Determination of Speech Feature Extraction Methods
In the pre - experimental stage, we compared different speech feature extraction methods, including mel - spectrogram features, chromaticity features, and contrast features. We tested the TESS dataset and found that the accuracy of the mel - spectrogram feature on the test set reached 99.82%2(a), significantly better than the chromaticity feature (97.86%)2(b) and contrast feature (87.41%)2(c). Based on these results, the mel - spectrogram feature showed stronger recognition ability and robustness, so we chose the mel - spectrogram as the optimal feature extraction method and widely used it in the training and evaluation of all models in subsequent experiments.
We trained five models (ResNet, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet), all with a similar number of parameters. During training, mel - spectrogram features were used and trained on the IEMOCAP dataset (60% training set, 40% test set). All models achieved nearly 100% accuracy on the training set, but performed differently on the test set. Specifically, the test set accuracy of ResNet was 41.60%3(a), MobileNetV3 was 31.94%3(b), ShuffleNetV2 was 34.13%3(c), SqueezeNet was 43.46%3(d), and GhostNet was 33.40%3(e). These results show that although the models performed well on the training set, they had weak generalization ability on the test set, reflecting the overfitting phenomenon, which provides a basis for the subsequent use of the Conformal Prediction framework for model evaluation.
(a) ResNet training and testing Loss-Acc
@@ 4.2 Determination of Speech Feature Extraction Methods
4.3 Empirical Results of the ECR metric
We used SCP and Risk Control frameworks to evaluate the coverage performance of five models on the IEMOCAP dataset, and applied these models to the TESS dataset for testing. On the IEMOCAP and TESS datasets, both frameworks provided reliable coverage guarantees and showed that the two frameworks had good generalization capabilities across different datasetsFigs. 4 and 5.
@@ 4.4 Uncertainty Estimation of the Model through the APSS metric
Table 1: The prediction efficiency of five IEMOCAP-trained models was quantified through Split Conformal Prediction and risk control, measuring average prediction set size (±SD) across two evaluation corpora (label counts specified parenthetically).
@@
0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Method: Split Conformal Prediction
IEMOCAP(6) Resnet $4.025_{\pm 0.049}$ $3.106_{\pm 0.050}$ $2.319_{\pm 0.044}$ $1.755_{\pm 0.040}$ $1.297_{\pm 0.030}$ $0.941_{\pm 0.025}$ $0.621_{\pm 0.023}$ $0.362_{\pm 0.019}$ $0.160_{\pm 0.011}$
Mobilenetv3 $4.570_{\pm 0.062}$ $3.705_{\pm 0.063}$ $2.948_{\pm 0.052}$ $2.346_{\pm 0.064}$ $1.818_{\pm 0.040}$ $1.344_{\pm 0.035}$ $0.922_{\pm 0.026}$ $0.529_{\pm 0.027}$ $0.217_{\pm 0.015}$
Shufflenetv2 $4.468_{\pm 0.065}$ $3.534_{\pm 0.048}$ $2.800_{\pm 0.053}$ $2.236_{\pm 0.047}$ $1.720_{\pm 0.041}$ $1.252_{\pm 0.028}$ $0.832_{\pm 0.032}$ $0.494_{\pm 0.022}$ $0.211_{\pm 0.016}$
Squeezenet $3.975_{\pm 0.072}$ $2.989_{\pm 0.051}$ $2.230_{\pm 0.048}$ $1.659_{\pm 0.044}$ $1.237_{\pm 0.031}$ $0.868_{\pm 0.021}$ $0.576_{\pm 0.025}$ $0.342_{\pm 0.017}$ $0.147_{\pm 0.011}$
Ghostnet $4.657_{\pm 0.064}$ $3.698_{\pm 0.065}$ $2.973_{\pm 0.052}$ $2.319_{\pm 0.052}$ $1.761_{\pm 0.037}$ $1.276_{\pm 0.028}$ $0.855_{\pm 0.027}$ $0.519_{\pm 0.020}$ $0.235_{\pm 0.016}$
TESS(4) Resnet $4.000_{\pm 0.000}$ $2.893_{\pm 0.059}$ $2.551_{\pm 0.035}$ $2.230_{\pm 0.041}$ $1.894_{\pm 0.044}$ $1.580_{\pm 0.040}$ $1.225_{\pm 0.029}$ $0.663_{\pm 0.120}$ $0.191_{\pm 0.017}$
Mobilenetv3 $4.000_{\pm 0.000}$ $4.000_{\pm 0.000}$ $4.000_{\pm 0.000}$ $2.210_{\pm 0.182}$ $1.846_{\pm 0.050}$ $1.526_{\pm 0.039}$ $1.161_{\pm 0.047}$ $0.662_{\pm 0.056}$ $0.281_{\pm 0.008}$
Shufflenetv2 $4.000_{\pm 0.000}$ $3.106_{\pm 0.558}$ $2.490_{\pm 0.051}$ $2.161_{\pm 0.047}$ $1.819_{\pm 0.042}$ $1.480_{\pm 0.038}$ $1.125_{\pm 0.034}$ $0.626_{\pm 0.047}$ $0.199_{\pm 0.022}$
Squeezenet $3.621_{\pm 0.051}$ $3.190_{\pm 0.055}$ $2.751_{\pm 0.060}$ $2.364_{\pm 0.064}$ $1.897_{\pm 0.056}$ $1.483_{\pm 0.059}$ $0.977_{\pm 0.045}$ $0.535_{\pm 0.042}$ $0.222_{\pm 0.018}$
Ghostnet $4.000_{\pm 0.000}$ $4.000_{\pm 0.000}$ $2.439_{\pm 0.026}$ $2.176_{\pm 0.043}$ $1.873_{\pm 0.040}$ $1.562_{\pm 0.052}$ $1.216_{\pm 0.035}$ $0.695_{\pm 0.059}$ $0.255_{\pm 0.024}$
Method: Risk Control
IEMOCAP(6) Resnet $4.017_{\pm 0.061}$ $3.091_{\pm 0.059}$ $2.306_{\pm 0.041}$ $1.759_{\pm 0.042}$ $1.294_{\pm 0.029}$ $0.935_{\pm 0.027}$ $0.619_{\pm 0.022}$ $0.361_{\pm 0.017}$ $0.155_{\pm 0.012}$
Mobilenetv3 $4.574_{\pm 0.060}$ $3.691_{\pm 0.058}$ $2.951_{\pm 0.049}$ $2.333_{\pm 0.043}$ $1.828_{\pm 0.042}$ $1.340_{\pm 0.042}$ $0.922_{\pm 0.027}$ $0.532_{\pm 0.025}$ $0.215_{\pm 0.015}$
Shufflenetv2 $4.464_{\pm 0.064}$ $3.553_{\pm 0.054}$ $2.801_{\pm 0.054}$ $2.230_{\pm 0.048}$ $1.713_{\pm 0.045}$ $1.240_{\pm 0.032}$ $0.823_{\pm 0.029}$ $0.495_{\pm 0.024}$ $0.210_{\pm 0.016}$
Squeezenet $3.964_{\pm 0.061}$ $2.979_{\pm 0.062}$ $2.217_{\pm 0.048}$ $1.666_{\pm 0.037}$ $1.234_{\pm 0.032}$ $0.868_{\pm 0.026}$ $0.569_{\pm 0.025}$ $0.344_{\pm 0.018}$ $0.149_{\pm 0.013}$
Ghostnet $4.649_{\pm 0.059}$ $3.707_{\pm 0.071}$ $2.979_{\pm 0.056}$ $2.313_{\pm 0.049}$ $1.764_{\pm 0.037}$ $1.271_{\pm 0.033}$ $0.859_{\pm 0.030}$ $0.518_{\pm 0.021}$ $0.235_{\pm 0.015}$
TESS(4) Resnet $4.000_{\pm 0.000}$ $2.887_{\pm 0.042}$ $2.555_{\pm 0.042}$ $2.219_{\pm 0.040}$ $1.904_{\pm 0.039}$ $1.581_{\pm 0.043}$ $1.223_{\pm 0.037}$ $0.666_{\pm 0.115}$ $0.188_{\pm 0.017}$
Mobilenetv3 $4.000_{\pm 0.000}$ $4.000_{\pm 0.000}$ $4.000_{\pm 0.000}$ $2.214_{\pm 0.051}$ $1.849_{\pm 0.041}$ $1.513_{\pm 0.044}$ $1.158_{\pm 0.043}$ $0.649_{\pm 0.061}$ $0.280_{\pm 0.009}$
Shufflenetv2 $4.000_{\pm 0.000}$ $2.840_{\pm 0.047}$ $2.494_{\pm 0.048}$ $2.153_{\pm 0.039}$ $1.814_{\pm 0.040}$ $1.477_{\pm 0.039}$ $1.126_{\pm 0.033}$ $0.611_{\pm 0.044}$ $0.196_{\pm 0.023}$
Squeezenet $3.613_{\pm 0.046}$ $3.192_{\pm 0.051}$ $2.751_{\pm 0.052}$ $2.359_{\pm 0.058}$ $1.921_{\pm 0.065}$ $1.477_{\pm 0.063}$ $0.978_{\pm 0.042}$ $0.530_{\pm 0.033}$ $0.221_{\pm 0.021}$
Ghostnet $4.000_{\pm 0.000}$ $3.975_{\pm 0.174}$ $2.486_{\pm 0.051}$ $2.174_{\pm 0.050}$ $1.865_{\pm 0.046}$ $1.556_{\pm 0.052}$ $1.215_{\pm 0.031}$ $0.676_{\pm 0.065}$ $0.249_{\pm 0.024}$
@@
We found a negative correlation between the average prediction set size (APSS) and the risk level. Specifically, as the risk level increases, the APSS value decreases significantly1, indicating that the model’s prediction results at high risk levels are more certain and the size of the prediction set is optimized. In this way, APSS is able to effectively evaluate the uncertainty of the classification model on the test set. Larger APSS values indicate that the model has higher prediction uncertainty for certain categories, while smaller APSS values mean that the model has higher confidence in its predictions. Therefore, APSS, as a measure of model uncertainty, demonstrates its potential as a promising benchmark in evaluating the robustness of classification models and provides a valuable reference for further optimizing models.
@@ 4.5 Ablation Studies
The experimental results show that although the calibration set size changes significantly (from 10% to 90%), this method always ensures the empirical coverage of the test set satisfies the theoretical lower bound $1-\alpha$ ($\alpha=0.2$)6 by dynamically adjusting the confidence threshold $q$. Theoretical analysis and empirical results jointly prove that this method exhibits strong robustness to data partition ratios. The constructed prediction sets provide statistically rigorous reliability guarantees for speech emotion recognition tasks, effectively addressing the confidence calibration failure problem in traditional methods caused by data distribution shifts.
4.6 Relax the exchangeability based on small batch martingale process
In order to verify the effectiveness of the proposed martingale method and the rationality of the assumption of local interchangeability in small batch scenarios, this paper adopts a cross-dataset experimental design, focusing on the actual scenario where inter-batch data does not meet global
(a) SCP method
(b) martingale method
Figure 7: Coverage guarantee based on martingale method
interchangeability. The calibration set is selected from the IEMOCAP speech emotion dataset, and the test set is the TESS dataset. There are significant differences in the speaker, emotional expression mode and acoustic environment, forming a typical domain offset problem, which directly leads to inconsistent data distribution between batches (that is, the inter-batch does not meet the interchangeability). The experiment divides the data into "local batches", each batch contains 5 calibration data and 1 test data, strengthening the local correlation of internal data by limiting the batch size - although there may be significant distribution differences between batches (such as grouping characteristics of different time periods and different scenarios), the data within a single batch can approximately meet "local interchangeability" because it comes from a closer local distribution.
As shown in the figure7, traditional methods experience significant coverage bias when data does not satisfy exchangeability, as the global exchangeability assumption fails, resulting in empirical coverage below the target value. This method uses the local exchangeability assumption and martingale value screening, and when the distribution difference between batches is significant, the empirical coverage rate and target value are still achieved, verifying the validity of the relaxation assumption that "interchangeability can not be met between batches".$M_{t}$ significantly affects the quality of the conformal set at batch t. As the martingale approaches zero, the conformal set becomes larger and less informative. The figure shows the changing path of $M_{t}$8. It is worth noting that this method ensures coverage but does not control the size of the conformal set; stronger models generally produce smaller sets.
Figure 8: Supermartingale Analysis
This framework does not require global data to obey the same distribution, significantly relaxing the prerequisites for traditional conformal prediction, and is suitable for complex scenarios where there is a distribution offset between batches.
@@ 5 Conclusion
This study proposes a solution based on CP and risk control framework to solve the problem of unreliable prediction of traditional neural networks in speech emotion recognition. The results show that traditional point prediction methods have significant confidence bias due to factors such as data distribution differences, model overfitting, and difficulty in capturing complex emotional features. By using the SCP framework and risk control framework, this study successfully transforms single-point predictions into set predictions with coverage guarantees, ensuring that the true label is included in the prediction set with a probability of at least $1-\alpha$ under a specified risk level $\alpha$. In addition, cross-dataset validation proves the robustness of the framework under different data distributions. The study also finds that APSS is negatively correlated with the risk level, indicating that when the risk level is high, the model’s prediction set is usually smaller and the model’s prediction uncertainty is lower. Based on this finding, we propose a metric for evaluating the uncertainty of classification models on the test set. At the same time, our mini-batch martingale value conformal prediction theory extends the applicability of CP to non-exchangeable environments, solving a key flaw in existing uncertainty quantification methods. By adopting the local exchangeability assumption within the mini-batch and constructing a non-negative test martingale, our method maintains coverage guarantees even under non-IID data distributions.